The last soldier of an 80 m long marching troops runs from the end to its front,and then it runs back to the end with same speed.During this,marching troops covers a distance of 150 m.The distance covered by the soldier is , a). 310 b). 250 c). 230 d). 160

Grade:12th pass

1 Answers

RANADIP TAH
7 years ago
Let v1 be the speed of last soldier & v2 be the speed of the marching troop. So relative velocity of soldier in forward motion is v1-v2 & relative vel. in backward motion is v1+v2.=> time(t) = distance / velocitySo, 150/v2 = 80/(v1-v2) + 80/(v1+v2) = time(t)=> v2= 150/t=> putting v2= 150/ t in the above equation we get,=> v1t^2 -160v1t - 22500=0=> v1t^2 -250v1t + 90v1t -22500=0=> (v1t -250)(v1t +90)=0 Now we get ,Distance covered by soldier = v1 * t = 250m (b) [answer]